OpenStudy (anonymous):

what is x^2-64 factored out to be

OpenStudy (anonymous):

I think it should be (x+8)(x-8)

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

Yay!(:

OpenStudy (anonymous):

thnk guys im like so out of it right now and I .. just thank you

hero (hero):

lolz

OpenStudy (anonymous):

Your welcome(:

hero (hero):

Also, just in case you didn't know, a^2 - b^2 = (a + b)(a - b) That's called Difference of Squares

OpenStudy (anonymous):

yeah thats what its call i forgot thank you

OpenStudy (anonymous):

i have one more tho

OpenStudy (anonymous):

what is it?

hero (hero):

This time, it's probably going to be the last question in the book. Something super hard, lol

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

Lol I might know, what is it? I'm an eighth grader taking 10th grade math, so who knows lol

OpenStudy (anonymous):

\[\frac{ 15x+5 }{ x ^{2} -1}\]

OpenStudy (anonymous):

I don't remember how to solve this one, sorry :(

OpenStudy (anonymous):

http://www.youtube.com/watch?v=w2YVIPAsoeA

OpenStudy (anonymous):

but thats not one the answers

hero (hero):

Did you post the question correctly? Double check.

OpenStudy (anonymous):

o no sorry on the numerator it is 15x+15 thank you

hero (hero):

\[\frac{15x + 15}{x^2-1}=\frac{15(x + 1)}{(x+1)(x-1)} = \frac{15\cancel{(x+1)}}{\cancel{(x+1)}(x-1)}\]

OpenStudy (anonymous):

thank you